OpenStudy (anonymous):

find the equation of the line containing the given pair of points. (-6,-6) (2,2)

OpenStudy (anonymous):

If you plot the graph, you will see that m=1 (45º). (m is the slope...) so using the formula : (y-yo)=m(x-xo) and replace the value of point (2,2) for example.. we'll have y-2=1(x-2) y=x

OpenStudy (anonymous):

Did you get?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok! have a good study!